What is the net present value (NPV) of your proposed expansion into the Canada? Assume that the cash flows after year 0 occur at the end of each year. The required rate of return is 17.5%. (Round to nearest penny) Year 0 cash flow = -730,000 Year 1 cash flow = -120,000 Year 2 cash flow = 400,000 Year 3 cash flow = 490,000 Year 4 cash flow = 540,000 Year 5 cash flow = 440,000
Answer #1

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=400,000/1.175^2+490,000/1.175^3+540,000/1.175^4+440,000/1.175^5

=$1071528.71

Present value of outflows=730,000+120,000/1.175

=832127.66

NPV=Present value of inflows-Present value of outflows

=1071528.71-832127.66

=$239401.05(Approx).